OpenStudy (zepp):

Integration problem: \[\large \int 7x^4(e^x)^5 dx\]I just don't know what to substitute :( I did one step though... \[\large 7 \int x^4(e^x)^5 dx\]Factored out the 7! :D Help please!

OpenStudy (calculusfunctions):

Is that \[e ^{x ^{5}}\] or \[(e ^{x})^{5}=e ^{5x}\] As soon as you clear that up, I can help you?

OpenStudy (zepp):

The whole thing to the 5th power

OpenStudy (calculusfunctions):

So you're saying the it's the second case??

OpenStudy (zepp):

Yes

OpenStudy (calculusfunctions):

OK! then you must use integration by parts. Do you know that or would you like to see a detailed solution?

OpenStudy (zepp):

Didn't learn Integration by parts yet. :/

OpenStudy (anonymous):

Apply tabular method for Integration by Part will be much easier!

OpenStudy (calculusfunctions):

Integration by parts:\[\int\limits_{}^{}udv = uv -\int\limits_{}^{vdu}\] Do you know this. If not then tell me and I'll show you. Or as Chlorophyll said, a tabular method works just as well.

OpenStudy (calculusfunctions):

Sorry, that's \[\int\limits_{}^{}vdu\]

OpenStudy (zepp):

Haven't seen this yet, I guess I'll wait a bit for this kind of exercices, thank you everybody anyway!

OpenStudy (calculusfunctions):

OK, if you haven't learned integration by parts, have you learned the tabular method?

OpenStudy (zepp):

I'm still in Calculus 1 stuffs

OpenStudy (zepp):

MIT's problem sets probably mixed Calc 1 and Calc 2 stuffs together D:

OpenStudy (calculusfunctions):

Well those are the two methods. Is it possible you weren't in class that day? Yes it is also possible that Calc 1 & 2 were combined. That's besides the point. Which method would you like to learn from me?

OpenStudy (zepp):

I don't take Calculus at school, I'm doing this on my own haha, I'll skip this one for now, and I gotta go, thank you @calculusfunctions for your help anyways!